OpenStudy (anonymous):

The area of a square garden is 200 m^2. how long is the diagonal?

OpenStudy (anonymous):

area of square = side^2 side = root(area) = root(200m^2) = 10root2 m using pythagoras theorem (diagonal)^2 = (10root2)^2 + (10root2)^2 = 200 + 200 diagonal = root400 = 20 m